Which of the following expressions shows the number of 8 character passwords that can be formed using letters and digits if the password must begin with a letter?

Answer by Theo(13342)

the number of letters in the alphabet is equal to 26 (a to z)
the number of numbers in a 1 digit number is equal to 10 (0 to 9)

if the first character has to be a letter, then you have 26 possible choices for that first character.
if the other characters can be a letter or a number, than you have 36 possible choices for each of the other 7 characters.

the number of possible choices is therefore 26 * 36^7.

if you allow for capital and small letters, then the number of characters in the alphabet becomes 52.

if those are not any of your choices, let me know what the choices are and i'll pick one for you if i can figure out what they're looking for.
the total number of choices then becomes 52 * 62^7

26 * 36^7 = 2.037... * 10^12

52 * 62^7 = 1.831... * 10^14

you can use your calculator to get the whole number if you need that.

when the calculator shows E14, that means * 10^14.